40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 244 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  4710   Fri May 13 01:42:35 2011 kiwamuUpdateLSCready for Schnupp asymmetry measurement

[Valera / Kiwamu]

 We are able to lock each arm smoothly.  It is ready for the Schnupp asymmetry measurement.

( to be done )

 - Manual D-phase adjustment of the AS55 channel.

 - A script to adjust the D-phase.

 - Trigger logic for the boost filters.

 - Modification of some old alignment scripts to adopt them to the new LSC channels

  4709   Fri May 13 00:39:53 2011 valeraUpdateASSc1ass update

Here the status of the dither alignment or c1ass:

- Both pitch and yaw centering on ETMY/ITMY were closed simultatenously with ugf of ~1/30 Hz.

- I made a medm screen with beam positions as measured by the dither system.The snapshot is attached. There are visual perimeter alarms (red box around the display) to warn about arm power being low or the dither lines not being on. The screen has a pull down menu with 4 scripts:

. assUp - sets up the gains, phases and matricies for the dither system (both the spot centering and the input beam alignment)

. assOn - turns on the dithers and servo - just the Y-arm centering part at the moment

. assOff - turns off the servo and dither lines

. assDitherOn - turns on the dither lines but does not turn on the servo

- All scripts are in scripts/ASS and the medm screen is in medm/c1ass/master/

 

Still to do:

- Commission the input beam and X-arm servos

- Make scripts for X-arm

Attachment 1: c1assqpds.jpg
c1assqpds.jpg
  4708   Thu May 12 23:50:10 2011 SureshUpdateRF SystemPOY55_Demod board Hardware change completed

The Demod board with S. No. 022 (being used earlier as REFL11) has been modified.  It now has SCLF-65 as its input LP filter on the PD input line and a PQW-2-90 power splitter.  The unit functioning okay (I and Q signals are 90 deg apart.

The loss of Q output was traced to a possible loose solder joint and we now have both the I and Q signals after resoldering all components in the vicinity of U7 (Ref Schematic of D990511)

There is a strong oscillation around 350Hz present on I and Q signals of both REFL55_Demod and POY55_Demod.  Don't know the source. 

We have run out of power splitters to continue with the Demod board modification. We do not currently have an AS11_Demod board.  All the others are in place and ready for the I<->Q phase angle measurement.

In summary we now have the following Demod boards in place:

[ REFL11, POY11, REFL55, AS55, POY55, POY22, POY110]_Demod

 

  4707   Thu May 12 23:41:47 2011 ranaUpdateCDSMEDM Snapshots not working

Looks like 2 different MEDM Snapshot functiions (at least) are broken.

The regular update of the screens here as well as the usual "Update Snapshot" and view "previous snapshot" button on all of the auto-generated screens.

Also, how do we add the snapshot button to the custom made screens?

  4706   Thu May 12 23:12:40 2011 kiwamuUpdateCDSc1lsc crashed

This is my third time to crash a real-time machine. This time I crashed c1lsc.

I physically rebooted c1lsc machine by pushing the power button and it came back and now running fine.

 

(what I did)

The story is almost the same as the last two times (1st time, 2nd time).

I edited c1lsc.ini file using daqconfig and then shutdown daqd running fb.

Some indicators for c1lsc on the C1_FE_STATUS screen became red. So I hit the 'DAQ reload' button on the C1LSC_GDS_TP screen.

Then c1lsc died and didn't respond to ping.

  4705   Thu May 12 22:54:20 2011 ranaUpdateDAQInput Beam Naming change (no more IP)

 We decided to rename the Input Beam channels (while keeping temporary backwards compatible aliases) as:

C1:ASC-IB_POS_X, C1:ASC-IB_POS_Y, C1:ASC-IB_ANG_SUM, etc.

  4704   Thu May 12 18:21:19 2011 kiwamuUpdateIOOwelcome back IP_POS

IP_POS is back. 

I reconnected the cable to an interface card : D030238-A which has been labeled as "IP POS".

The card currently resides on the third crate rack from the top at the very right side in 1Y2. Also a rear side connection was modified a little bit.

I will clean up some cables because I tried some cables to see which was which.

Hopefully I will make a simple wiring diagram such that we will never forget the connections.

Quote from #4700

Quote from #4799

A cable for IP_POS has been disconnected at the LSC rack, 1Y2. Due to it currently no IP_POS signal shows up on the digital side.

I will check how it used to be and reconnect it.

 I found the disconnected cable, but I do not see the interface board at the LSC rack

  4703   Thu May 12 16:50:22 2011 SureshUpdateRF Systeminstallation of RF splitters in Demod boards

We have no plan to change the AS11 PD. I was referring to the AS11 Demod board which currently has the "Demodulator Preamp" circuit installed as a piggy back. In future I will append "_Demod" when I am referring to a demod board, to avoid confusion.

Quote:

Quote:

Rana also advised that we must use the boards which have the piggy-back amplifiers on those signals which are most useful.  We referred to Alberto's thesis and chose POY55 (MICH and SRCL), REFL11(PRCL)  and AS55 (DARM) as the most useful signals.  We currently have these amps on AS11, REFL11 and AS55.  We need to convert either AS11 or REFL11 into a POY55.  Since we need to troubleshoot REFL11, I thought we might as well modify that and in the process also fix its Q output.  So I renamed AS11 as REFL11 and will convert the old REFL11 into POY55 tomorrow. 

 I think we should leave them as is; the AS11 was made by taking into account the SB levels at the AS port and should not become REFL11. We should instead convert one of the old 25 or 33 MHz diodes into a POY55.

 

  4702   Thu May 12 10:27:02 2011 ranaUpdateRF Systeminstallation of RF splitters in Demod boards

Quote:

Rana also advised that we must use the boards which have the piggy-back amplifiers on those signals which are most useful.  We referred to Alberto's thesis and chose POY55 (MICH and SRCL), REFL11(PRCL)  and AS55 (DARM) as the most useful signals.  We currently have these amps on AS11, REFL11 and AS55.  We need to convert either AS11 or REFL11 into a POY55.  Since we need to troubleshoot REFL11, I thought we might as well modify that and in the process also fix its Q output.  So I renamed AS11 as REFL11 and will convert the old REFL11 into POY55 tomorrow. 

 I think we should leave them as is; the AS11 was made by taking into account the SB levels at the AS port and should not become REFL11. We should instead convert one of the old 25 or 33 MHz diodes into a POY55.

  4701   Thu May 12 09:00:23 2011 steveUpdateGeneraloptical table covers must be used

I found ITMX-ISCT and BS-ISCT ( IFO sensing-control table)  not covered this morning. Please do not leave tables open.

  4700   Thu May 12 08:54:11 2011 steveUpdateIOOIP_POS cable found

Quote:

A cable for IP_POS has been disconnected at the LSC rack, 1Y2. Due to it currently no IP_POS signal shows up on the digital side.

It looks like we disconnected the cable together with some unused cables when we were cleaning up the wiring of the LSC rack.

The cable, a shielded flat-cable, is supposed to send DC power to the QPD and send the signals from the QPD to an interface board on the LSC rack.

I will check how it used to be and reconnect it.

 I found the disconnected cable, but I do not see the interface board at the LSC rack

  4699   Thu May 12 05:25:26 2011 kiwamuUpdateIOOIP_POS disconnected

A cable for IP_POS has been disconnected at the LSC rack, 1Y2. Due to it currently no IP_POS signal shows up on the digital side.

It looks like we disconnected the cable together with some unused cables when we were cleaning up the wiring of the LSC rack.

The cable, a shielded flat-cable, is supposed to send DC power to the QPD and send the signals from the QPD to an interface board on the LSC rack.

I will check how it used to be and reconnect it.

  4698   Thu May 12 02:31:07 2011 SureshUpdateRF Systeminstallation of RF splitters in Demod boards

[Jamie, Koji, Suresh]

We replaced splitters in several demod boards as per the table given below:

 

Demod boards in which splitter has been replaced on 11May2011
Demod Board S. No. Power Splitter Frequency range of splitter (MHz) Phase unbalance from datasheet (deg) Amplitude Unbalance from datasheet (dB)
REFL55 118 PQW-2-90 30 to 90 90.14 0.92
AS11 021 PSCQ-2-51W 5 to 50 87.49 0.1
POY11 119 PSCQ-2-32 3.2 to 32 87.58 0.05
POY22 020 PSCQ-2-32 3.2 to 32 90.26 0.02
POY110 120 PSCQ--120 80 to 120 90.88 0.58

 

While doing a rough check of the boards I noticed that the REFL11 demod board had no signal on the Q output.

Rana also advised that we must use the boards which have the piggy-back amplifiers on those signals which are most useful.  We referred to Alberto's thesis and chose POY55 (MICH and SRCL), REFL11(PRCL)  and AS55 (DARM) as the most useful signals.  We currently have these amps on AS11, REFL11 and AS55.  We need to convert either AS11 or REFL11 into a POY55.  Since we need to troubleshoot REFL11, I thought we might as well modify that and in the process also fix its Q output.  So I renamed AS11 as REFL11 and will convert the old REFL11 into POY55 tomorrow.

Power splitter of different types have different pin-outs.  The way we mount a splitter depends on which type we are using.   I will detail the mounting scheme in a separate elog tomorrow.

 

 

 

  4697   Thu May 12 00:59:45 2011 ryan, ranaUpdatePSLReturn of the PSL temperature box

The PSL temperature box has returned to service, with some circuit modifications. The 1k resistors on all the temp. sensor inputs (R3, R4, R7, R8, R12, R12) were changed to 0 Ohm. Also, the 10k resistors R26, R28, R29, and R30 were changed to 10.2k metal film. The DCC document will be updated shortly. There is now an offset in the MINCOMEAS channel compared to the others, which will be corrected in the morning after looking at the overnight trend.

  4696   Wed May 11 23:02:52 2011 valeraUpdateASSDither angular stabilizitaion system update

This is what was done in past two days:

- The ETMY and ITMY pitch and yaw dofs are modulated at 40, 44, 42, 46 Hz respectively (oscillator A=30). The c1ass lockin numbers are 12, 14, 27, 29.

- The NAS55I signal is demodulated at the above frequencies. The demodulated I/Q signal phase is set to shift all signal into I-phase. The lockin inputs are bandpassed around respective frequency f with butter("Bandpass",2,f-0.5,f+0.5). The demod signals are then additionally low passed with butter ("Lowpass",4,0.5) so the servo ugf has to be below 0.5 Hz. The servo filter is p:z 0.0001:0.1.

- The ETMY demodulated signal is fed back to ITMY and visa versa.

- With the above 2x2 servo running we moved the input beam PZTs by hand to follow the cavity.

- At the end we offloaded the servo control signals to the SUS biases again by hand.

- The beam spot centering was estimated by unbalancing the ETMY/ITMY pitch/yaw coil combinations intentionally by 5%, which produces 1.3 mm shift of the node, and comparing the response to the residual signals.

- The dof set up currently is: ETMY pitch lockin 12 -> dof2, ITMY pitch lockin 14 -> dof4, ETMY yaw lockin 27 -> dof7, ITMY yaw lockin 29 -> dof9

- The next step is to demodulate the TRY(X) and servo the input beam PZTs

  4695   Wed May 11 22:54:53 2011 JenneUpdateTreasureETMY trans QPD installed

I put the ETMY trans QPD in. 

The ETMY trans beam was already going toward the TRY DC PD, and a CCD camera.  I put a beam splitter in that beam (reducing the power to TRY and the CCD by 50%), and sent my picked-off beam to the ETMY QPD.  Since there is a lens in this path to focus the beam onto TRY and the camera, I put the QPD ~the same distance from the lens as the camera.  Due to space requirements (because of all the green stuff on the table now), I had to put a Y1 turning mirror between the beam splitter and the QPD.  The beam is aligned onto the PD, although the signal isn't super strong.  When the PD is blocked, the sum is ~(-92 counts).  When the beam is on the PD, the sum is ~(-78 counts). 

  4694   Wed May 11 22:52:55 2011 kiwamuUpdateLSCstatus update and plan

Rana forced me to write this entry for summary because he didn't come to the 40m meeting today.

Status update :

    Interferometer Input Beam alignment with the PZTs.

     60 % done. The rest of the 40 % is to make the procedures automated.

     The beam spots on ITMY and ETMY are centered within ~1 mm accuracy.

     PRM, BS, ITMX, & ITMY actuator calibrations and PRC/MICH error calibrations

    Ongoing: First we will do it by hand, then some scripts will be made for the calibraion and resultant noise budget.

     F2A Suspension filter calculations.

      ETMY and ITMY are done. Need volunteer for ETMX, ITMX, BS, PRM, & SRM !!

      Bounce-Roll notch filters

      Leo is working on it. 25% complete...

     DC signal from RFPDs

      The RFPDs have a local SMA DC output as well as a DC output from their PD Interface cards in the LSC rack. We have hooked up some of the PD Interface DC outputs to the LSC ADCs but not tested??

Next steps:

  Installation of a temporary (Thorlabs) DCPD on either POY to see the intra-cavity power in the PRC. It would be ridiculous to put detectors on POX or POP since they're still clipped.

  D-phase and amplitude imbalance adjustment of the demod baords.

          make a script which uses pynds.

  Alignment of the full interferomter, starting from the X arm

  Loss measurements for the arms

  Schnupp asymmetry measurement

  

  4693   Wed May 11 21:58:30 2011 ranaUpdateSUSETMX_OPLEV : He-Ne replaced

Quote:

Quote:

I found that a He-Ne laser which has been used for ETMX_OPLEV was NOT giving the light.

Since I didn't find the switch key for it I have no idea if the laser is simply off or dead.

 The dead laser was replaced by new JDSU 1104P of  2.6mW. The return beam is big ~5 mm diameter of 0.3 mW,  1400 counts

 Whenever replacing any Oplev laser, please also put into the ELOG when it was installed so that we have an electronic record of the laser lifetime.

  4692   Wed May 11 17:20:24 2011 kiwamuConfigurationIOOloop diabled on PZT2

[Valera / Kiwamu]

The pointing of the incident beam to the interferometer has been jumping frequently.

Due to this jump the lock of the Y arm didn't stay for more than 2 min.

We turned off the strain gauge loop of PZT2-YAW and PZT2-PITCH, then the spot motion became solid and the Y arm locking became much more robust.

  4691   Wed May 11 17:10:04 2011 kiwamuConfigurationElectronicsfixed : MC3 LL PD has no signal

[Valera / Kiwamu]

It was because of a loose connection. Pushing the connector solved the issue.

looseconnectionMC3.png

We really have to think about making reliable connections and strain reliefs.

Quote from #4685

Yesterday we found that MC3 OSEM LL PD did not have a sensible signal - the readback was close to zero and it was making MC move around. I disabled the PD LL so that the damping is done with just three face plus side PDs. There still no signal from MC3 LL PD today. It needs debugging.

 

  4690   Wed May 11 16:04:36 2011 Larisa ThorneConfigurationElectronicsJenne-Seismometer LPF project

The schematic for the seismometer box from this last time  has been updated...

 

Koji was helpful for coming up with a general diagram for the voltage buffer amplifier, which has now been added to the configuration pictured below.

The only thing that remains now before I try to plot it with Eagle/LISO is to pick an op amp to use for the voltage buffer itself. Someone suggested THS4131 for that (upon Googling, it hit as a "high speed, low noise, fully-differential I/O amplifier"). It looks good, but is it the best option?

Attachment 1: STS2diagram2.png
STS2diagram2.png
  4689   Wed May 11 13:54:32 2011 steveUpdateLSCLSC power supplies

Aux- rack _1Y2 is just behind 1Y3  It contains Kepco Analoge DC power supplies for +- 5, 15 and 24V

Placing these power supplies away from the LSC rack was an effort to minimize pick up from them.

 

Attachment 1: P1070679.JPG
P1070679.JPG
  4688   Wed May 11 11:44:52 2011 kiwamuUpdateSUSf2p filters installed on ITMY

New f2p filters were installed on ITMY this morning. The statistical error of the coil gain setttigs are about 0.6 % at high frequency.

NEXT : PRM, SRM, BS, ITMX and ETMX.

 

f2p_itmy.png

 Pendulum mode = 0.988 Hz, Q = 1

C1:SUS-ITMY_ULCOIL_GAIN = 1.02482

C1:SUS-ITMY_URCOIL_GAIN = -1.06831

C1:SUS-ITMY_LLCOIL_GAIN = -0.996671

C1:SUS-ITMY_LRCOIL_GAIN = 0.91079

UL: fz = 1.014824 Hz, Q = 1.027150

UR: fz = 0.975038 Hz, Q = 0.98688

LL: fz = 1.000229 Hz, Q = 1.012378

LR: fz = 0.972688 Hz, Q = 0.946116

Quote from #4682

New f2p filters were installed on ETMY. 

The coil imbalances are now about 0.8% at high frequency (i.e. above the resonant freq of the pendulum mode)

  4687   Wed May 11 11:24:59 2011 steveUpdateSUSSRM-oplev qpd is back

Quote:

The SRM qpd cable was removed from the BS-table. It's path was changed from 1x4 to ITMY-table following the inner cable tray.

 Laser diode oplev SRM is working. Qpd matrix values were reset like others.

In 0.44mW, returning 0.1mW,  -500 counts.

Attachment 1: srm-prm_oplevs.jpg
srm-prm_oplevs.jpg
  4686   Wed May 11 11:05:45 2011 steveUpdateSUSETMX_OPLEV : He-Ne replaced

Quote:

I found that a He-Ne laser which has been used for ETMX_OPLEV was NOT giving the light.

Since I didn't find the switch key for it I have no idea if the laser is simply off or dead.

 The dead laser was replaced by new JDSU 1103P of  2.6mW. The return beam is big ~5 mm diameter of 0.3 mW,  1400 counts

Attachment 1: P1070672.JPG
P1070672.JPG
  4685   Wed May 11 10:49:16 2011 valeraConfigurationElectronicsMC3 LL PD has no signal

Yesterday we found that MC3 OSEM LL PD did not have a sensible signal - the readback was close to zero and it was making MC move around. I disabled the PD LL so that the damping is done with just three face plus side PDs. There still no signal from MC3 LL PD today. It needs debugging.

  4684   Wed May 11 00:23:21 2011 kiwamuUpdateASCY arm and beam pointing alignment

[Jamie / Valera / Kiwamu]

 The incident beam pointing was improved a lot by using C1ASS realtime code.

Some more details will be posted later. The below is the list of the highlights today.

 - The Y arm cavity was aligned to have good beam centering on the mirrors.

 - The input PZTs were also aligned to the aligned Y arm by hand.

 - Automation of the  Y arm alignment using C1ASS_LOCKIN got partially functional with two loops closed. C1ASS correctly servos the centering on ETMY

 - The amount of the off-centering on ITMY and ETMY look roughly within 1 mm.

 - As a result the intracavity power got bigger by a factor of about 3.5

  4683   Tue May 10 23:14:01 2011 kiwamuUpdateSUSETMX_OPLEV : He-Ne laser dead ?

I found that a He-Ne laser which has been used for ETMX_OPLEV was NOT giving the light.

Since I didn't find the switch key for it I have no idea if the laser is simply off or dead.

  4682   Tue May 10 22:56:17 2011 kiwamuUpdateSUSf2p filters installed on ETMY

New f2p filters were installed on ETMY. 

The statistical error of the coil gain settings are now about 0.8% at high frequency (i.e. above the resonant freq of the pendulum mode)

 What I did :

 -   measured and corrected the coil imbalances on ETMY using a script called F2P_LOCKIN.py

 -   made the new f2p filters based on the measurements and installed them.

  Next step :

 -  do the same adjustment for all the suspensions including PRM, SRM, BS, ITMs and ETMs

 


(Notes on F2P_LOCKIN.py)

 F2P_LOCKIN.py is a script that I've made in python. This is basically the same as the old script, f2pratio, but uses the realtime LOCKINs instead of ezcademods.

The script automatically measures the coil imbalances on an optics of interest by driving the local LOCKIN oscillators.

In the first step the script automatically balances the coil gains at high frequency (8.5Hz).

In the next step it gives some coefficients, which basically represent the coil imbalances at low frequency (0.01Hz)

Then with those coefficients one will be able to design the f2p filters.

It is not well polished yet, so I will spend some more times to make it user-friendly and readable.

Example usage : F2P_LCKIN.py -o ETMX

It currently resides in /cvs/cds/rtcds/caltech/c1/scriptss/SUS/

 

(new f2p filters)

The plot below shows the new f2p filters. Note that they are already installed.

f2p_etmy.png

Pendulum mode = 0.982 Hz (according to the wiki)
Q of the pendulum mode = 1 (to avoid ringing of IIR filters)
C1:SUS-ETMY_ULCOIL_GAIN = -0.793435
C1:SUS-ETMY_URCOIL_GAIN = 1.16877
C1:SUS-ETMY_LLCOIL_GAIN = 1.25028
C1:SUS-ETMY_LRCOIL_GAIN = -0.730704
UL_fz = 0.965125
UR_fz = 1.029517
LL_fz = 0.934231
LR_fz = 0.996562
UL_Qz = 0.982816
UR_Qz = 1.048388
LL_Qz = 0.951356
LR_Qz = 1.014829
  4681   Tue May 10 20:57:05 2011 ranaUpdatePSLPSL reference cavity temperature box modifications : after 24 hours still OK

Quote:

I re-installed the box (@ ~8:15) after reflowing some of the solder joints. I will observe it over night and then remove the 1K resistors. Attached is a 8 hour minute-trend.

 I compared this 24 hour trend with the one from this day exactly one year ago. Seems the same, so now I can make the resistor change.

Attachment 1: Untitled.png
Untitled.png
  4680   Tue May 10 16:45:19 2011 josephbUpdateCDSc1ass now receiving AS55I from c1lsc

[Valera, Joe]

We added a cdsPCIx_SHMEM connection between the c1lsc and c1ass models.  This connection is called C1:LSC-ASS_AS55I, and sends the normalized AS55I data to Lockin 11 of the c1ass model.

In addition, in order to get the c1ass model to compile, we had to place all the non-IO parts inside a subsystem block, which we called ASS, and gave the top_names tag to.

The c1lsc and c1ass models were rebuilt, the frame builder restarted, and the models restarted.

  4679   Tue May 10 16:42:49 2011 josephbUpdateCDSNew upconversion model (c1uct)

[Ryan, Joe]

Ryan added the c1uct (upconversion tester) model to the c1ioo machine.   It is DCU_ID 32, CPU 6.  The relevant wiki page has been updated. It has been added to /diskless/root/etc/rtsystab on the fb machine and automatically comes up when the c1ioo computer is turned on. 

Joe still needs to add its status to the status screen.

It is soft linked from:

/opt/rtcds/caltech/c1/userapps/trunk/CDS/c1/models/c1uct.mdl

Ryan will expand upon its uses later.

  4678   Tue May 10 11:14:44 2011 steveUpdatePSLPSL-QPD_ANG qpd optimized

Valera and I placed F 572.7 mm lens ~15 cm away  from the ang  qpd (in the same mount with ND filter) so that two qpds see different combination of ang and pos motion - there was no lenses prior to this change. The beam diameter is reduced to ~half .

Attachment 1: qpdangl.jpg
qpdangl.jpg
  4677   Tue May 10 10:06:23 2011 josephbUpdatePSLFixed channel names for PSL QPDs, fixed saturation, changed signs

I added calculation entries to the /cvs/cds/caltech/target/c1iool0/c1ioo.db file which are named C1:IOO-QPD_*_*, as the channels were originally named.  These calculation channels have the identical data to the C1:PSL-QPD_*_* channels.  I then added the channels to the C0EDCU.ini file, so as to once again have continuity for the channels, in addition to having the newer, better named channels.

The c1iool0 machine ("telnet c1iool0", "reboot") and the framebuilder process ("telnet fb 8087", "shutdown") were both restarted after these changes.

These channels were brought up in dataviewer and compared.  The approriate channels were identical.

Quote:

[Valera, Joe]

Software Changes:

First we changed all the C1:IOO-QPD_*_* channels to C1:PSL-QPD_*_* channels in the /cvs/cds/caltech/target/c1iool0/c1ioo.db file, as well as the /opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini file.  We then rebooted the frame builder via "telnet fb 8087" and then "shutdown".

This change breaks continuity for these channels prior to today.

 

  4676   Tue May 10 09:05:49 2011 KojiUpdateGeneralaLIGO BBPD / Green PD investigation

BBPD update

- The BBPD circuit has been constructed on the aLIGO BBPD board

- It still keeps 200MHz BW with FDD-100 Si PD for the 100V bias.

- The noise spectrum has been cleaned up a lot more. It shows the noise level of the 0.4mA shotnoise between 9-85MHz.
The noise at 160MHz is the noise level of the 1mA shotnoise.

Some of the noise peaks at around 97MHz came from the bias voltage.

What to do next

- Confirmation of the performance with the original aLIGO BB PD configuration.

- Notch filter for 9MHz (for aLIGO).

- Implementation of a power amplifier. (issues: power supply and heat removal)

Attachment 1: PD_response.pdf
PD_response.pdf PD_response.pdf PD_response.pdf
  4675   Tue May 10 01:39:41 2011 SureshUpdateRF SystemRF source troubles

Today after Steve and I finished the RF cabling work for the day, Kiwamu noticed that there were no RF signals to be seen.  The problem was traced to disconnected 11 and 55 MHz Demod lines from the RF source.  But reconnecting them did not restore the signals.  It turned out that one of the Heliax cables had a loose N-type connector at its end and it finally came off while we were tightening it into place.

We replaced the damaged heliax with another (we have two spare running from 1X2 IOO rack to the 1Y2 LSC rack.  The new cable is used to be the LO 33.  It seems to have a 1.5dB loss.  Have to check this again tomorrow.

In the mean time I noticed that the power output of the 55MHz Demod port of the source was less than about -12dB. So I opened the source to take a look and found that all the voltage stabilisers were supplying 15V.  Even those which were supposed to be supplying 24V.  This was traced to a mistake in wiring the external power supply.  The wires had been labeled wrongly and as a result the 18V input line was connected to 28V source and vice versa.

After fixing this problem I reassembled the source checked the power output on all the ports and found everything was functioning as expected.  However after installation once again the unit failed.  The blue light on the power supply was not lighting up when switched on.  Suspecting a power supply problem I opened the unit again and found that a weak solder joint on one of the RF amplifiers had come loose and had overloaded one of the 24V stabilisers.   We, found a spare and replaced it.  The unit has been reassembled and is functioning fine.  The output power levels are

11MHz Demod -- 6dBm

55MHz Demod -- 5.5 dBm

11MHz EOM --   24dBm

55MHz EOM -- 28dBm

The Marconi is serving as the 11MHz source.  The Wenzel 11MHz source is giving 13.3 dBm and is okay.  But it needs to be checked for its performance as it may have been exposed to higher than rated power supply levels.

The 29.5MHz source is giving 7dBm.  It is supposed to be giving 13dBm. 

The Laboratory DC power supplies currently used for both the RF source and Distribution boxes need to be replaced with rack mounted Sorensen power supplies available in the lab.

 

  4674   Tue May 10 00:44:52 2011 valeraUpdateIOOMC2 centering

Kiwamu, Koji, Valera

We centered the beam on MC2 in pitch by moving the MC1,2,3 in the following combination [-9,+3,-7]. This actuation vector mostly moves the spot on MC2 vertically. The attached plot shows the dither before and after the centering. We monitored the demodulated signals and saw the reduction of the MC2 pit response from -1.0 to -0.22 which corresponds to the beam spot position change from 9 to 2 mm. Thus all the spots on MC mirrors are within 2 mm of the center. We estimate based on the distance between the MC1-MC3 of 20 cm, the distance from the center between MC1 and MC3 to the end of the Faraday isolator of 80 cm, and the aperture of the FI of 12 mm, the maximum angle out of MC of 3/200 rad. Which implies the maximum differential spot motion of 3 mm not to be limited by the FI aperture.

Attachment 1: mc2centering.pdf
mc2centering.pdf
  4673   Tue May 10 00:31:28 2011 kiwamuUpdateASCIFO alignment plan

The alignment of the interferometer goes basically step by step.

Tuesday will be an alignment day.

  0. MC beam centering (it's done)

  1. F2P to balance the coils on every optics including BS, PRM, SRM, ITMs and ETMs (Kiwamu).

  2. A2L and then change the DC bias of ITMY and ETMY to get a perfect eigen axis (VF/Jamie).

  3. align input PZT mirrors (PZT1 and 2) to maximize the Y arm transmission (VF/Jamie).

  4. do the same things for X arm but using BS instead of the PZTs.

  5. Alignment of the central part.

  6. Make a script to automatically get those things done. 

  4672   Mon May 9 20:30:20 2011 ranaUpdatePSLPSL reference cavity temperature box modifications

I re-installed the box (@ ~8:15) after reflowing some of the solder joints. I will observe it over night and then remove the 1K resistors. Attached is a 8 hour minute-trend.

Attachment 1: Untitled.png
Untitled.png
  4671   Mon May 9 18:54:29 2011 kiwamuUpdateLSCtriggerd locking logic : screen prepared

I made a new medm screen for the triggering logics. Have fun.

I put a button on C1LSC.adl to invoke this screen.

trigger_medm.png

Quote from #4656

It will be quite useful for the Y arm locking, for instance we can do a triggered locking and the maximization of the intracavity power.

 

  4670   Mon May 9 17:23:25 2011 SureshUpdateRF SystemRF Cables near LSC Rack

[Steve, Suresh]

We started to clean up the RF cables (heliax and PD interface cables)  at the LSC rack.

We have pulled out all the RF cables from the small hole on the side-board close to floor.  Passing the cables through this hole makes some of the cables much too short for good strain relief.  So we removed the side panel on the vacuum tube side and are going to pass the cables into the rack from there at about waist height.  We now have plenty of cable lengths to tie them off to the rack at several points.

We have traced all the available Heliax cables and have attached blank tags to them.  We have allocated some cables to REFL11, REFL55 and AS55.  These are therefore back in working order.  We have also taken stock of the available PD interface cables.  They do not have consistent names on both ends of the cable and we will identify and label the ends tomorrow.

MC is locked.  The auto-locker works fine.

Handing over the system for night time interferometer work now.  Will continue with the cabling tomorrow.

 

  4669   Mon May 9 17:21:13 2011 JamieBureaucracyCDS!!!!REMEMBER TO COMMIT YOUR MODIFICATIONS!!!!

Also, remember to update the svn working copy before you begin doing any work, to make sure you're working on the most recent version:

cd /opt/rtcds/caltech/c1/userapps/trunk
svn update
  4668   Mon May 9 16:58:23 2011 JamieBureaucracyCDS!!!!REMEMBER TO COMMIT YOUR MODIFICATIONS!!!!

Whenever you modify any of the front-end code in /opt/rtds/caltech/c1/userapps, such as models, scripts or medm screens, REMEMBER TO COMMIT YOUR CHANGES, WITH A LOG MESSAGE!!!

cd /opt/rtcds/caltech/c1/userapps/trunk
svn commit path/to/modified/file

If you forget to commit things, they may be purged and you will loose your work.  ALWAYS COMMIT!

Things to note and watch out for:

  • When you commit, make sure you specify explicitly the path to the file that you are committing.  This repository is shared by all LIGO sites, including the sites (LHO and LLO) so you really want to make sure you're not committing anything that affects anything other than "c1" (ie. the 40m)
  • Only commit one file at a time, unless the changes to multiple files are tightly coupled.
  • Never commit without specifying the path to the file that you are committing.  Other people may be working on other files in the repository, and you don't want to commit their work with yours.
  • YOU MUST LEAVE A LOG MESSAGE.  Your log message should be specific to the file you have modified, and should be clear and verbose to explain exactly what you have done.
  • Use the "svn status" command to give you an overview of what is going on in the working directory.  This will tell you which files have been modified.  See "svn help status" for more information.

If you have questions ASK.  Don't force things.

  4667   Mon May 9 16:12:49 2011 JamieConfigurationCDScanonicalize paths to core and userapps

I have updated the /opt/rtcds paths to reflect the new specification of the CDS aLIGO code release procedures document.


Path to RTS/opt/rtcds/caltech/c1/core/release

This is where the advLigoRTS front-end code generator is checked out.  The "release" directory is a link to the svn branch from which we are currently running ("trunk" by default).

This used to be at /opt/rtcds/caltech/c1/core/advLigoRTS


Path to userapps: /opt/rtcds/caltech/c1/userapps/release

This is where the cds_user_apps code is checked out.  cds_user_apps is where all of the front-end models, medm screen, scripts, etc. will live.  The "release" directory is a link to the svn branch from which we are currently running ("trunk" by default).

This was most recently at /opt/rtcds/cds_user_apps

 

  4666   Mon May 9 15:21:36 2011 josephbUpdatePSLFixed channel names for PSL QPDs, fixed saturation, changed signs

[Valera, Joe]

Software Changes:

First we changed all the C1:IOO-QPD_*_* channels to C1:PSL-QPD_*_* channels in the /cvs/cds/caltech/target/c1iool0/c1ioo.db file, as well as the /opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini file.  We then rebooted the frame builder via "telnet fb 8087" and then "shutdown".

This change breaks continuity for these channels prior to today.

The C1:PSL-QPD_POS_HOR and C1:PSL-QPD_POS_VERT channels were found to be backwards as well.  So we modified the /cvs/cds/caltech/target/c1iool0/c1ioo.db file to switch them.

Lastly, we changed the ASLO and AOFF values for the C1:PSL-QPD_POS_SUM and the C1:PSL-QPD_ANG_SUM so as to provide positive numbers.  This was done by flipping the sign for each entry.

ASLO went from 0.004883 to -0.004883, and AOFF when from -10 to 10 for both channels.

Hardware Changes:

The C1:PSL-QPD_ANG_SUM channel had been saturated at -10V.  Valera reduced the power on the QPD to drop it to about 4V by placing an ND attenuator in the ANG QPD path.

  4665   Mon May 9 13:14:48 2011 josephbUpdateLSCC1:LSC-TRIG_MTRX : wrong matrix size

[Joe, Kiwamu]

There is a feature/bug of the RCG code that you can only have 1 receiving tag for every sending tag.  There were 5 tags which were being received by two tags each, for two different matrices.  Only the first tag was receiving, the second was apparently ignored.

This has been fixed temporarily by putting in direct lines in place of these 5 tags.

Quote:

I found that C1:LSC-TRIG_MTRX has a wrong matrix size. It needs to be fixed.

It is designed to have a 11x8 matrix in the simlink model file, but it's been compiled as a 6x8 matrix.

 

  4664   Mon May 9 12:33:40 2011 kiwamuUpdateLSCC1:LSC-TRIG_MTRX : wrong matrix size

I found that C1:LSC-TRIG_MTRX has a wrong matrix size. It needs to be fixed.

It is designed to have a 11x8 matrix in the simlink model file, but it's been compiled as a 6x8 matrix.

I recompiled c1lsc but it didn't fix the issue. Here below is the matrix statement in the C file :c1lsc.c. Indeed it is 6x8 matrix ....

// MuxMatrix:  LSC_TRIG_MTRX
for(ii=0;ii<8;ii++)
{
  lsc_trig_mtrx[ii] =
          pLocalEpics->lsc.LSC_TRIG_MTRX[ii][0] * lsc_imux_trigger[0] +
          pLocalEpics->lsc.LSC_TRIG_MTRX[ii][1] * lsc_imux_trigger[1] +
          pLocalEpics->lsc.LSC_TRIG_MTRX[ii][2] * lsc_imux_trigger[2] +
          pLocalEpics->lsc.LSC_TRIG_MTRX[ii][3] * lsc_imux_trigger[3] +
          pLocalEpics->lsc.LSC_TRIG_MTRX[ii][4] * lsc_imux_trigger[4] +
          pLocalEpics->lsc.LSC_TRIG_MTRX[ii][5] * lsc_imux_trigger[5];
 }

  4663   Mon May 9 09:37:51 2011 valeraUpdatePSLPSL and MC trends

The attached plot shows 7 day trends of the MC and PMC power levels, PSL QPDs, and temperature. The MC stayed locked for ~40 hours over the weekend. The temperature swings were somewhat smaller over the past couple of days but one should remember to turn the PSL HEPA down after working on the table. Steve turned the HEPA flow from 100% down to 20% on Thursday and posted the reminder signs on the PSL enclosure.

Attachment 1: pslmcdrift2.pdf
pslmcdrift2.pdf
  4662   Sun May 8 22:59:40 2011 ranaUpdatePSLPSL reference cavity temperature box modifications

I looked at the PSL temperature box. It started out as D980400-B-C. Then it was revised by Peter King as per the LHO mods E020247.

There are some more things to do to it to make it useful for us:

  1. R3, 4, 7, 8, 12, & 13 should be changed from 1k to 0 Ohms, I think. I cannot figure out their purpose.
  2. All resistors should be made metal-film. Right now, its kind of a mish-mash.
  3. The active filter U6B has a corner frequency of ~50 Hz. This seems not useful for keeping the 4116 DAC noise out of the temperature. We should lower this to ~30 mHz to take advantage of the stability of the LT1021 which was put in.

** Frank reminds me that we don't use the TIdal or VME external inputs anymore since we moved to the EPICS/Perl PID control. So all we have to do is make sure these inputs are hardware disabled/disconnected.

  4661   Sun May 8 17:29:01 2011 ranaUpdateIOOMC beam spot centering

It seems like the best option would be to make the MCASS just adjust the SUS biases and center the beams on the suspended optics. Is this not possible somehow?

ELOG V3.1.3-